OpenStudy (anonymous):

During a storewide sale, a TV which usually sells for $450 is on sale for $360. A Stereo which usually sells for $600 is on sale for $480. What would be the sale price of a VCR which usually sells for $500?

OpenStudy (ddcamp):

Find x, then find ? 480 = 600x 360 = 450x ? = 500x

OpenStudy (anonymous):

360/450=.9 480/600=.9 x/500=.9 Solve for x. x=400

OpenStudy (ddcamp):

@jeetster 480/600 & 360/450 = 0.8, but your final answer is still right.

OpenStudy (anonymous):

@DDCamp whoops, the problem with seeing the answer in your head before you see the steps your brain took.